# On Selecting Few-Shot Examples for LLM-based Code Vulnerability Detection

### v1

@@
Abstract.
Large language models (LLMs) have demonstrated impressive capabilities for many coding tasks, including summarization, translation, completion, and code generation. However, detecting code vulnerabilities remains a challenging task for LLMs. An effective way to improve LLM performance is in-context learning (ICL) – providing few-shot examples similar to the query, along with correct answers, can improve an LLM’s ability to generate correct solutions. However, choosing the few-shot examples appropriately is crucial to improving model performance. In this paper, we explore two criteria for choosing few-shot examples for ICL used in the code vulnerability detection task. The first criterion considers if the LLM (consistently) makes a mistake or not on a sample with the intuition that LLM performance on a sample is informative about its usefulness as a few-shot example. The other criterion considers similarity of the examples with the program under query and chooses few-shot examples based on the k-nearest neighbors to the given sample. We perform evaluations to determine the benefits of these criteria individually as well as under various combinations, using open-source models on multiple datasets.
LLMs, Few-shot examples, Code vulnerability detection
00footnotetext: Preprint. Under review.
1. Introduction
@@
Large language models (LLMs) have demonstrated impressive capabilities for coding tasks including detection of software defects and vulnerabilities (such as resource leaks, use-after-free of dynamic memory during program operation, and denial-of-service attacks), variable misuse detection (i.e., programmers used wrong variables), code summarization (i.e., represent a code segment with a single word), and code generation and code completion (Jiang et al., 2024; Zhou et al., 2025; Sun et al., 2025; Husein et al., 2025; Wan et al., 2024). In this paper we explore the applicability of LLMs for vulnerability detection. This is a particularly important application, as software vulnerabilities are prevalent in many software systems, posing serious risks such as compromising sensitive data and system failures. There are limited studies in previous work that show that state-of-the-art LLMs with few-shot-learning capabilities can achieve competitive results in detecting software vulnerabilities compared to previous machine learning techniques (Zheng et al., 2024; Lin and Mohaisen, 2025). However, the performance (measured in terms of precision, recall, F1 score) of these models remains low (Ding et al., 2024; Khare et al., 2025) preventing the deployment of these models in realistic settings.
@@
In this paper, we explore black-box, prompt-based methods for improving the performance of LLMs on vulnerability detection. Our idea is based on improving the efficacy of in-context learning (ICL)—the well-known phenomenon of LLM’s exhibiting improved performance when the provided prompt includes a set of examples (referred to as the few-shot set) that demonstrate the task to be performed—for the purpose of vulnerability detection. The impact of ICL on model performance is highly sensitive to the specific examples chosen and therefore selecting the appropriate few-shot examples is crucial to improving model performance (Liu et al., 2021; Min et al., 2022; Xu et al., 2023b). For vulnerability detection, the few-shot set comprises of programs along with the ground-truth label indicating whether they have a vulnerability or not.
@@
We propose new methods for choosing the example programs that comprise the few-shot set used in the code vulnerability detection task. Our first contribution is to present a pair of algorithmic primitives for choosing the few-shot examples. The first primitive, Learn-from-Mistakes (LFM), is based on the idea that LLM performance on a sample is informative about its usefulness as a few-shot example. Therefore, LFM first queries the model (multiple times) on a potential example and records if the model is consistently mistaken (or correct) on the example. An example is added to the few-shot set only if the the model consistently makes mistakes on it. The intuition is that the model does not know well how to reason about that example therefore adding it in the few-shot set will likely help rectify the model behavior, while a randomly chosen example may not be as useful. We also explore alternate versions of LFM where an example is added only if the model is consistently correct on it. The intuition here is that adding examples where the model is consistently correct can help reinforce the desired behavior and improve performance. The second primitive, Learn-from-Nearest-Neighbors (LFNN), is based on the intuition that the semantic similarity between an example and the program under test can be a helpful guide in deciding if the example should be added to the few-shot set. In particular, LFNN adds the nearest neighbors to the set. To compute semantic similarity, LFNN relies on code embedding models that map programs to embedding vectors. Similarity of vectors can be computed using metrics such as cosine similarity.
@@
Our second contribution is to propose three different methods for combining LFM and LFNN to yield few-shot sets that are both semantically similar to the program under test and provide a context summarizing the model’s past mistakes. Each combination yields a unique algorithm to construct the few-shot set for the vulnerability detection task and represents a specific trade-off between the two criteria—model performance and semantic similarity—for selecting examples.
@@
We evaluate our methods using two popular open-source models that are known to perform well on coding tasks, namely Qwen2.5-Coder-7B-Instruct and Gemma-3-4B-it, as well as a closed-source model, GPT-5-mini. We use four challenging datasets for our evaluation that include programs from a variety of languages such as C/C++, Python, and JavaScript. Our experimental results indicate that individual strategies, particularly LFNN, can significantly enhance the models’ baseline vulnerability detection capabilities. LFM introduces a strong inductive bias, consistently skewing predictions toward the positive class. However, our proposed combined methods demonstrate a more robust and balanced performance profile, effectively optimizing both accuracy and F1-score across the diverse range of models and datasets evaluated.
2. Related Work
@@
Vulnerability detection is a long standing challenge in software security. Over the years, a wide range of techniques have been developed to detect vulnerabilities, ranging from static program analysis (Lipp et al., 2022; Marques et al., 2025) to dynamic testing and fuzzing (Manès et al., 2021; Cassel et al., 2023, 2025), and from these traditional methods to more recent machine learning based approaches (Li et al., 2025a, 2018; Hin et al., 2022; Melicher et al., 2021).
Machine learning based approaches aim to detect vulnerabilities by learning patterns directly from data. Early work extracted hand-crafted features such as token frequencies, code complexity metrics, or dependency patterns, and trained classifiers to separate vulnerable from non-vulnerable code (Nam and Kim, 2015; Medeiros et al., 2020). More recent methods use representation learning to automatically encode code or abstract representations of code into vector spaces that possibly capture both syntax and semantics, enabling models to recognize patterns that are difficult to design manually. These approaches have shown promising results (Zhou et al., 2019; Li et al., 2021a; Hin et al., 2022; Steenhoek et al., 2024; Yang et al., 2024b), but their effectiveness is often constrained by dataset quality and the challenge of generalizing across diverse projects (Croft et al., 2023; Yadav and Wilson, 2024; Li et al., 2025b; Safdar et al., 2025).
@@
Recent attempts to apply large models (LM) to vulnerability detection generally fall into two broad approaches. The first line of work treats LMs as embedding models, extracting vector representations of code and then either training lightweight classifiers on top of these embeddings or fine tuning the LM itself for the downstream task (Hin et al., 2022; Hanif and Maffeis, 2022; Steenhoek et al., 2023). Many of these methods also incorporate a program’s abstract representations such as Abstract Syntax Trees (ASTs), Control Flow Graphs (CFGs), or data flow features to complement the raw code input (Steenhoek et al., 2024; Yang et al., 2024b; Zhou et al., 2019; Li et al., 2021b). Early research in this category has often focused on models like CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020), with more recent work exploring larger foundation models such as Qwen (Bai et al., 2023; Team, 2024; Yang et al., 2024a, 2025) and LLaMA (Roziere et al., 2023; Dubey et al., 2024; Touvron et al., 2023).
The second line of work leverages the generative capabilities of modern large language models (LLM) more directly. Instead of relying only on embeddings, these approaches test the ability of LLMs to reason about and classify vulnerabilities through code understanding and natural language generation. Some studies evaluate models in their pretrained form, while others fine tune the models to improve task performance (Zhou et al., 2025). Several recent efforts reflect this trend. Du et al. (2024a) introduce VulLLM, a multi-task LLM framework that integrates vulnerability interpretation and data augmentation to significantly improve code vulnerability detection. Farr et al. (2025) leverage few-shot prompting to enhance out-of-the-box LLMs for vulnerability detection. Du et al. (2024b) proposes knowledge-level retrieval-augmented generation (RAG) for code vulnerability detection and reports sizable gains over baselines.
@@
In-context learning (ICL) (Brown et al., 2020) enables a model to adapt to a new task by conditioning on a small number of input output examples provided in the prompt. These examples serve as implicit supervision, enabling the model to infer the task format and desired output style without expensive fine tuning. ICL has been widely studied in natural language and code related tasks, where it has been shown to substantially improve model performance over zero-shot prompting (Brown et al., 2020; Ramesh et al., 2025; Wei et al., 2022; Bansal et al., 2019).
A central question in ICL is how to select the few examples that are most useful for the model. Selection strategies can range from simple random sampling to more sophisticated methods that consider similarity between the test input and candidate examples, prior evidence about LLM behavior, or task specific heuristics (Liu et al., 2021; Min et al., 2022; Jiang et al., 2020; Rubin et al., 2022; Xu et al., 2023b). Liu et al. (2021) retrieve examples that are semantically-similar to a test sample to formulate its corresponding prompt. Rubin et al. (2022) train an efficient dense retriever to select training examples as prompts at test time. Xu et al. (2023b) employ in-context learning to create expert profiles that condition LLM responses. In our setting, we investigate new strategies for selecting few shot examples and study how they influence the performance of LLMs on vulnerability detection.
@@
Prompt optimization broadly refers to techniques that improve how tasks are presented to large language models so that they yield more reliable outputs (Pryzant et al., 2023; Sabbatella et al., 2024). Beyond manual design, recent work has begun to explore automated prompt optimization. DSPy (Khattab et al., 2023) introduces a declarative framework that compiles language model pipelines into self-improving programs. GEPA (Agrawal et al., 2025) uses reflective natural-language feedback and evolutionary search to iteratively refine prompts. Maestro (Wang et al., 2025) extends this line of work to multi-agent settings by jointly optimizing node configurations and the structure of agent graphs to better mitigate structural failure modes. These advances connect closely to the problem of example selection in ICL, as the choice and arrangement of input–output demonstrations influences model performance (Lu et al., 2022; Guo et al., 2024; Agarwal et al., 2024; Bhope et al., 2025). In our work, we treat the construction of the few-shot example set as a form of prompt optimization and study how different selection strategies affect the performance of LLMs for vulnerability detection.
3. Two Algorithms: LFM and LFNN
@@
3.1. Learn-from-Mistakes (LFM)
@@ 17 $\mathcal{S}_{I}\leftarrow\mathcal{S}_{I}\cap\mathcal{S}_{Ii}$;
21 if stst then
@@ 27 $\mathcal{S}\leftarrow\text{Uniformly draw }n\text{ samples from }S_{opt}$;
Algorithm 1 chooses few-shot examples based on the intuition that the correctness of the LLM response on an example is informative about its usefulness as a few-shot example. The algorithm makes a linear scan over a labeled dataset—for each sample in the dataset, it queries an LLM asking it to predict if the sample program has a vulnerability or not. This information is used to construct a few-shot set from this dataset (which we refer to as the training dataset). In its default version, the algorithm, which we call Learn-from-Mistakes (LFM), operates under the assumption that the examples on which the LLM makes mistakes are more informative than random examples and should be added to the few-shot set. It iteratively updates the few-shot set based on incorrect predictions. However, the algorithm can be configured to run under various other settings that we describe next.
@@
Finally, the few-shot set $\mathcal{S}$ to be returned by the algorithm is computed (lines 16-21). If the flag $st$ is set to TRUE, the examples in the initial set $\mathcal{S}_{init}$ are included in the set $\mathcal{S}$. The remaining samples in the set $\mathcal{S}$ are chosen from the appropriate sets $\mathcal{S}_{C}$, $\mathcal{S}_{I}$, and $\mathcal{S}_{G}$ (denoted by $\mathcal{S}_{opt}$) as dictated by the value of the optopt input (lines 18 and 21).
3.2. Learn-from-Nearest-Neighbors (LFNN)
@@ 13 $\textbf{return }\mathcal{NN}_{x}$;
Algorithm 2, which we call the Learn-from-Nearest-Neighbors (LFNN) algorithm, chooses few-shot examples based on the intuition that the program samples most similar to the program under query are the most helpful in improving LLM performance. Although similar ideas have been proposed in the context of other applications (Xu et al., 2023a), in this work, we use this idea to improve the performance of LLMs for the vulnerability detection task.
@@ 18 $\textbf{return }LFM(f_{VD},\mathcal{S}_{LFNN},\mathcal{S}_{LFM},n_{1},st,k,opt)$
In contrast to methods 1 and 2 that both invoke LFM and LFNN just once, method 3 (lines 9-12) invokes LFM twice. This method first invokes LFM (line 10) in a manner similar to method 1, generating a few-shot set $\mathcal{S}_{LFM}$ of size $n_{2}$. This first call to LFM is query-agnostic and therefore, only needs to made once. Next, the method invokes LFNN (line 11), again in a manner similar to method 1 and generates a set of size $n_{3}$. As usual, the call to LFNN is query-specific and needs to be repeated for each query. Next, and unlike the other methods, LFM is invoked a second time. For this invocation, instead of using $D$ as the dataset, the few-shot set $\mathcal{S}_{LFNN}$ computed by LFNN is used as the dataset. This enables inclusion of only those examples in the final few-shot set that are most similar to the query while also accounting for the model correctness on these examples. Moreover, the second call to LFM uses the few-shot set $\mathcal{S}_{LFM}$ computed on line 10 as the initial set. The intuition here is that initializing LFM with these examples can make LFM aware of the examples on which the model makes a mistake (or is correct, depending on the optopt parameter) and thus, enable LFM to pick more effective examples for the final few-shot set. Note that the second call to LFM is also query-specific.
5. Experiments

### v2

# On the Difficulty of Selecting Few-Shot Examples for Effective LLM-based Vulnerability Detection

@@
Abstract
Large language models (LLMs) have demonstrated impressive capabilities across a wide range of coding tasks, including summarization, translation, completion, and code generation. Despite these advances, detecting code vulnerabilities remains a challenging problem for LLMs. In-context learning (ICL) has emerged as an effective mechanism for improving model performance by providing a small number of labeled examples within the prompt. Prior work has shown, however, that the effectiveness of ICL depends critically on how these few-shot examples are selected. In this paper, we study two intuitive criteria for selecting few-shot examples for ICL in the context of code vulnerability detection. The first criterion leverages model behavior by prioritizing samples on which the LLM consistently makes mistakes, motivated by the intuition that such samples can expose and correct systematic model weaknesses. The second criterion selects examples based on semantic similarity to the query program, using k-nearest-neighbor retrieval to identify relevant contexts.
We conduct extensive evaluations using open-source LLMs and datasets spanning multiple programming languages. Our results show that for Python and JavaScript, careful selection of few-shot examples can lead to measurable performance improvements in vulnerability detection. In contrast, for C and C++ programs, few-shot example selection has limited impact, suggesting that more powerful but also more expensive approaches, such as re-training or fine-tuning, may be required to substantially improve model performance.
publicationid: pubid: Workshop on LLM Assisted Security and Trust Exploration (LAST-X) 2026 27 February 2026, San Diego, CA, USA ISBN 978-1-970672-05-3 https://dx.doi.org/10.14722/last-x.2026.23025 www.ndss-symposium.org 11footnotetext: These authors contributed equally to this work.
I Introduction
@@
Large language models (LLMs) have demonstrated impressive capabilities for coding tasks including detection of software defects and vulnerabilities (such as resource leaks, use-after-free of dynamic memory during program operation, and denial-of-service attacks), variable misuse detection (i.e., programmers used wrong variables), code summarization (i.e., represent a code segment with a single word), and code generation and code completion [1, 2, 3, 4, 5]. In this paper we explore the applicability of LLMs for vulnerability detection. This is a particularly important application, as software vulnerabilities are prevalent in many software systems, posing serious risks such as compromising sensitive data and system failures. There are limited studies in previous work that show that state-of-the-art LLMs with few-shot-learning capabilities can achieve competitive results in detecting software vulnerabilities compared to previous machine learning techniques [6, 7]. However, the performance (measured in terms of precision, recall, F1 score) of these models remains low [8, 9] preventing the deployment of these models in realistic settings.
@@
In this paper, we explore black-box, prompt-based methods for improving the performance of LLMs on vulnerability detection. We seek to evaluate techniques for improving the efficacy of in-context learning (ICL)—the well-known phenomenon of LLM’s exhibiting improved performance when the provided prompt includes a set of examples (referred to as the few-shot set) that demonstrate the task to be performed—for the purpose of vulnerability detection. Previous works indicate that the impact of ICL on model performance is highly sensitive to the specific examples chosen  [10, 11, 12]. For vulnerability detection, the few-shot set comprises programs along with the ground-truth label indicating whether they have a vulnerability or not.
@@
We explore methods for choosing the example programs that comprise the few-shot set used in the code vulnerability detection task. In particular, we explore two natural algorithms for choosing the few-shot examples. The first algorithm, Learn-from-Mistakes (LFM), is based on the idea that LLM performance on a sample is informative about its usefulness as a few-shot example. Therefore, LFM first queries the model (multiple times) on a potential example and records if the model is consistently mistaken (or correct) on the example. An example is added to the few-shot set only if the model consistently makes mistakes on it. The intuition is that the model does not know well how to reason about that example therefore adding it in the few-shot set will likely help rectify the model behavior, while a randomly chosen example may not be as useful. We also explore alternate versions of LFM where an example is added only if the model is consistently correct on it. The intuition here is that adding examples where the model is consistently correct can help reinforce the desired behavior and improve performance. The second algorithm, Learn-from-Nearest-Neighbors (LFNN), is based on the intuition that the semantic similarity between an example and the program under test can be a helpful guide in deciding if the example should be added to the few-shot set. In particular, LFNN adds the nearest neighbors to the set. To compute semantic similarity, LFNN relies on code embedding models that map programs to embedding vectors. Similarity of vectors can be computed using metrics such as cosine similarity.
@@
We further explore different ways for combining LFM and LFNN to yield few-shot sets that are both semantically similar to the program under test and provide a context summarizing the model’s past mistakes. Each combination yields a unique algorithm to construct the few-shot set for the vulnerability detection task and represents a specific trade-off between the two criteria—model performance and semantic similarity—for selecting examples.
@@
We evaluate these methods using two popular open-source models that are known to perform well on coding tasks, namely Qwen2.5-Coder-7B-Instruct and Gemma-3-4B-it, as well as a closed-source model, GPT-5-mini. We use four challenging datasets for the evaluation that include programs from a variety of languages such as C/C++, Python, and JavaScript.
Our experimental results show that while certain standalone strategies, particularly LFNN, can improve baseline vulnerability detection performance, their effectiveness varies across models and datasets. LFM introduces a strong inductive bias that tends to skew predictions toward the positive class, which often limits its applicability. In contrast, the combined methods provide more stable and generally more balanced performance, although their benefits are not uniform. Depending on the underlying model and dataset, these combinations may yield meaningful gains in both accuracy and F1-score, or only modest or no improvements, reflecting the heterogeneous nature of the vulnerability detection task.
II Related Work
@@
Vulnerability detection is a long-standing challenge in software security. Over the years, a wide range of techniques have been developed to detect vulnerabilities, ranging from static program analysis [13, 14] to dynamic testing and fuzzing [15, 16, 17], and from these traditional methods to more recent machine-learning-based approaches [18, 19, 20, 21].
Machine-learning-based approaches aim to detect vulnerabilities by learning patterns directly from data. Early work extracted hand-crafted features such as token frequencies, code complexity metrics, or dependency patterns, and trained classifiers to separate vulnerable from non-vulnerable code [22, 23]. More recent methods use representation learning to automatically encode code or abstract representations of code into vector spaces that possibly capture both syntax and semantics, enabling models to recognize patterns that are difficult to design manually. These approaches have shown promising results [24, 25, 20, 26, 27], but their effectiveness is often constrained by dataset quality and the challenge of generalizing across diverse projects [28, 29, 30, 31].
@@
Recent attempts to apply large models (LM) to vulnerability detection generally fall into two broad approaches. The first line of work treats LMs as embedding models, extracting vector representations of code and then either training lightweight classifiers on top of these embeddings or fine tuning the LM itself for the downstream task [20, 32, 33]. Many of these methods also incorporate a program’s abstract representations such as Abstract Syntax Trees (ASTs), Control Flow Graphs (CFGs), or data flow features to complement the raw code input [26, 27, 24, 34]. Early research in this category has often focused on models like CodeBERT [35] and GraphCodeBERT [36], with more recent work exploring larger foundation models such as Qwen [37, 38, 39, 40] and LLaMA [41, 42, 43].
The second line of work leverages the generative capabilities of modern large language models (LLM) more directly. Instead of relying only on embeddings, these approaches test the ability of LLMs to reason about and classify vulnerabilities through code understanding and natural language generation. Some studies evaluate models in their pretrained form, while others fine tune the models to improve task performance [2]. Several recent efforts reflect this trend. [44] introduce VulLLM, a multi-task LLM framework that integrates vulnerability interpretation and data augmentation to significantly improve code vulnerability detection. [45] leverage few-shot prompting to enhance out-of-the-box LLMs for vulnerability detection. [46] proposes knowledge-level retrieval-augmented generation (RAG) for code vulnerability detection and reports sizable gains over baselines.
@@
In-context learning (ICL) [47] enables a model to adapt to a new task by conditioning on a small number of input output examples provided in the prompt. These examples serve as implicit supervision, enabling the model to infer the task format and desired output style without expensive fine tuning. ICL has been widely studied in natural language and code related tasks, where it has been shown to substantially improve model performance over zero-shot prompting [47, 48, 49, 50].
A central question in ICL is how to select the few examples that are most useful for the model. Selection strategies can range from simple random sampling to more sophisticated methods that consider similarity between the test input and candidate examples, prior evidence about LLM behavior, or task specific heuristics [10, 11, 51, 52, 12]. [10] retrieve examples that are semantically-similar to a test sample to formulate its corresponding prompt. [52] train an efficient dense retriever to select training examples as prompts at test time. [12] employ in-context learning to create expert profiles that condition LLM responses. In our setting, we investigate new strategies for selecting few-shot examples and study how they influence the performance of LLMs on vulnerability detection.
@@
Prompt optimization broadly refers to techniques that improve how tasks are presented to large language models so that they yield more reliable outputs [53, 54]. Beyond manual design, recent work has begun to explore automated prompt optimization. DSPy [55] introduces a declarative framework that compiles language model pipelines into self-improving programs. GEPA [56] uses reflective natural-language feedback and evolutionary search to iteratively refine prompts. Maestro [57] extends this line of work to multi-agent settings by jointly optimizing node configurations and the structure of agent graphs to better mitigate structural failure modes. These advances connect closely to the problem of example selection in ICL, as the choice and arrangement of input–output demonstrations influences model performance [58, 59, 60, 61]. In our work, we treat the construction of the few-shot example set as a form of prompt optimization and study how different selection strategies affect the performance of LLMs for vulnerability detection.
III Two Algorithms: LFM and LFNN
@@
III-A Learn-from-Mistakes (LFM)
@@ 17 $\mathcal{S}_{I}\leftarrow\mathcal{S}_{I}\cap\mathcal{S}_{Ii}$;
21 $\mathcal{S}_{opt}\leftarrow\mathcal{S}_{I}$;
22if $opt=C$ then
23 $\mathcal{S}_{opt}\leftarrow\mathcal{S}_{C}$
24
25else if $opt=G$ then
26 $\mathcal{S}_{opt}\leftarrow\mathcal{S}_{G}$
27
28
29if stst then
@@ 35 $\mathcal{S}\leftarrow\text{Uniformly draw }n\text{ samples from }S_{opt}$;
Algorithm 1 chooses few-shot examples based on the intuition that the correctness of the LLM response on an example is informative about its usefulness as a few-shot example. The algorithm makes a linear scan over a labeled dataset—for each sample in the dataset, it queries an LLM asking it to predict if the sample program has a vulnerability or not. This information is used to construct a few-shot set from this dataset (which we refer to as the training dataset). In its default version, the algorithm, which we call Learn-from-Mistakes (LFM), operates under the assumption that the examples on which the LLM makes mistakes are more informative than random examples and should be added to the few-shot set. It iteratively updates the few-shot set based on incorrect predictions. However, the algorithm can be configured to run under various other settings that we describe next. Although some variants may also select correct or gray examples, we hold onto the name Learn-from-Mistakes because the selection process is defined through mistake-based scanning. Mistakes serve as the primary signal for identifying informative examples.
@@
Finally, the few-shot set $\mathcal{S}$ to be returned by the algorithm is computed (lines 21-26). If the flag $st$ is set to TRUE, the examples in the initial set $\mathcal{S}_{init}$ are included in the set $\mathcal{S}$. The remaining samples in the set $\mathcal{S}$ are chosen from the appropriate sets $\mathcal{S}_{C}$, $\mathcal{S}_{I}$, and $\mathcal{S}_{G}$ (denoted by $\mathcal{S}_{opt}$) as dictated by the value of the optopt input (lines 23 and 26).
III-B Learn-from-Nearest-Neighbors (LFNN)
@@ 13 $\textbf{return }\mathcal{NN}_{x}$;
Algorithm 2, which we call the Learn-from-Nearest-Neighbors (LFNN) algorithm, chooses few-shot examples based on the intuition that the program samples most similar to the program under query are the most helpful in improving LLM performance. Although similar ideas have been proposed in the context of other applications [62], in this work, we use this idea to improve the performance of LLMs for the vulnerability detection task.
@@ 18 $\textbf{return }LFM(f_{VD},\mathcal{S}_{LFNN},\mathcal{S}_{LFM},n_{1},st,k,opt)$
In contrast to methods 1 and 2 that both invoke LFM and LFNN just once, method 3 (lines 9-12) invokes LFM twice. This method first invokes LFM (line 10) in a manner similar to method 1, generating a few-shot set $\mathcal{S}_{LFM}$ of size $n_{2}$. This first call to LFM is query-agnostic and therefore, only needs to be made once. Next, the method invokes LFNN (line 11), again in a manner similar to method 1 and generates a set of size $n_{3}$. As usual, the call to LFNN is query-specific and needs to be repeated for each query. Next, and unlike the other methods, LFM is invoked a second time. For this invocation, instead of using $D$ as the dataset, the few-shot set $\mathcal{S}_{LFNN}$ computed by LFNN is used as the dataset. This enables inclusion of only those examples in the final few-shot set that are most similar to the query while also accounting for the model correctness on these examples. Moreover, the second call to LFM uses the few-shot set $\mathcal{S}_{LFM}$ computed on line 10 as the initial set. The intuition here is that initializing LFM with these examples can make LFM aware of the examples on which the model makes a mistake (or is correct, depending on the optopt parameter) and thus, enable LFM to pick more effective examples for the final few-shot set. Note that the second call to LFM is also query-specific.
V Experiments